OpenStudy (ivanzud):

Urgent:Action Required::: The temperature of a chemical reaction ranges between 40 degrees Celsius and 180 degrees Celsius. The temperature is at its lowest point when t = 0, and the reaction completes 1 cycle during a 12-hour period. What is a cosine function that models this reaction?

OpenStudy (ivanzud):

help!

OpenStudy (ivanzud):

is it f(t) = 70 cos 12t + 110?

OpenStudy (anonymous):

C.

OpenStudy (cbarredo1):

C

OpenStudy (ivanzud):

why C?

OpenStudy (anonymous):

the amplitude A of your function is: \[\Large A = \frac{{180 - 40}}{2} = ...?\]

OpenStudy (ivanzud):

70

OpenStudy (cbarredo1):

Yes

OpenStudy (anonymous):

Corrcet

OpenStudy (ivanzud):

but c has -70

OpenStudy (anonymous):

since the period of our function is 12 hours, so the shape og our function has to be this: \[\Large \cos \left( {\frac{{2\pi }}{{12}}t} \right)\]

OpenStudy (ivanzud):

but why -70 in the front?

OpenStudy (anonymous):

since at t=0, the value is 40, then we have this drawing:  now, if I set t=0 into the first option I get: f(0)= 180 so the first option is not the right option

OpenStudy (ivanzud):

ok so it is C.

OpenStudy (cbarredo1):

Yes C

OpenStudy (anonymous):

please substitute t=0 into the third option, what do you get? f(0)=...? -70 cos + 110? we have: \[\Large - 70\left( {\cos 0} \right) + 110 = - 70 \times 1 + 110 = 40\] So da answer is C. U WELCOME IF ANY OF YALL NEED HELP IN MATH JUST ASK ME IM A LITTLE GOOD IN MATH k PEACE OUT SON